OpenStudy (anonymous):

in a lift that is accelerating upwards at 2m/s^2 a spring balance shows the apparent weight of a body to be 2.5 kg wt. what would be the reading if the lift were at 2m/s^2

OpenStudy (anonymous):

@hartnn can you help me?

OpenStudy (anonymous):

@Hero can you help me??

OpenStudy (kropot72):

Both values of acceleration in your question are the same. If that is correct then the answer would be 2.5 kg wt.

OpenStudy (anonymous):

oops.. i bad.. the question is wrong.. here is the proper one!

OpenStudy (anonymous):

in a lift that is accelerating downwards at 1m/s^2 a spring balance shows the apparent weight of a body to be 2.5 kg wt. what would be the reading if the lift were at 2m/s^2?

OpenStudy (kropot72):

Let g = 9.81 m/s^2 Let m be the mass of the body 2.5 = 9.81m - (m * 1) = 8.81m m = 2.5/8.81 Let x = reading at 2m/s^2 x = 9.81m - (m * 2) = 7.81m Now you can substitute for m and find x.